# The Euclidean distance degree of curves: from rational to line multiview varieties

### v1

@@ 4 Conclusion and future directions
In this article, we considered one-dimensional multiview varieties. We used topological arguments to give the first theoretical results for ED degrees of multiview varieties anchored at a Schubert variety. One of the key features of our proof is that each genericity condition is the complement of a union of hypersurfaces whose defining equations only rely on a pair of cameras or a single camera. This resulted in a surprising Section 2 to determine the ED degrees of multiview varieties where cameras have a prescribed structure. This is relevant for families of “dual cameras” found in resectioning [4] as well as calibrated cameras [14, 19] used in practice. We leave for future work the task of of generalizing these ideas to higher dimensional multiview varieties. This becomes more difficult because the camera centers may intersect the world variety. We also leave for future work the problem of determining the ED degree of line multiview varieties using a different embedding into the Grassmannian [16, 17, 5].

### v2

@@ 5 Conclusion and future directions
In this article, we considered one-dimensional multiview varieties. We used topological arguments to give the first theoretical results for ED degrees of multiview varieties anchored at a Schubert variety. One of the key features of our proof is that each genericity condition is the complement of a union of hypersurfaces whose defining equations only rely on a single camera or a camera pair. This resulted in a surprising Section 2 to determine the ED degrees of multiview varieties where cameras have a prescribed structure. This is relevant for families of “dual cameras” found in resectioning [6] as well as calibrated cameras [21, 26] used in practice. A natural extension is to generalize our results on the ED degree of curves to higher-dimensional varieties. We leave for future work the task of combining the techniques in [27], which were used to prove (1.5), with the approaches we outlined in our proofs to give a generalization of our Section 2 for all parameterized anchored mutliview varieties. Another interesting problem is to determine the ED degree of line multiview varieties using alternative embeddings of the Grassmannian, such as those appearing in recent work on Grassmannian optimization [23, 24, 7]. A natural next step is to choose a representation of the embedded multiview variety that yields good properties for the bundle adjustment problem. For example, we can ask how the study of representations for 3D lines in projective space in [3, Section 3] generalizes to other multiview varieties. We hope that further developments along these lines will deepen the interplay between algebraic geometry, multilinear algebra, and computer vision.
Acknowledgments
We thank Timothy Duff for his comments on early versions of the article.
Data availability statement
The calculations supporting the results of this article can replicated using the open source software Macaulay2 [16] using the EuclideanDistanceDegree.m2 package. The code for these calculations is available at the GitHub repository.